OpenStudy (anonymous):

The area of a triangle is 124 square units. What would its new area be if its base was half as long, and the height was three times as long?

jimthompson5910 (jim_thompson5910):

A = (b*h)/2 124 = (b*h)/2 (1/2)*124 = ((b/2)*h)/2 ... replace b with b/2...this means we must multiply the left side by 1/2 62 = ((b/2)*h)/2 3*62 = ((b/2)*(3*h))/2 ... multiply both sides by 3 (to go from h to 3h) 186 = ((b/2)*(3*h))/2 So the new area is 186 square units

OpenStudy (anonymous):

:D Thank you for explaining! :D

jimthompson5910 (jim_thompson5910):

yw